The city fuel economy of a 2009 Toyota 4Runner 2WD 6 cylinder 4 L automatic 5-speed using regular gas is a normally distributed random variable with a range 21 MPG to 26 MPG.

(a)

Estimate the standard deviation using Method 3 (the Empirical Rule for a normal distribution). (Round your answer to 4 decimal places.)

  Standard deviation   
(b)

What sample size is needed to estimate the mean with 98 percent confidence and an error of ± 0.25 MPG? (Enter your answer as a whole number (no decimals). Use a z-value taken to 3 decimal places in your calculations.)

  Sample size   
Answer #1

sd deviahon yang method g± tand ar L The empi) ca sue normal dishbuhon] der, ation Cơ) -maximum valW-min, man Vale 6 σ= 26-212 2 | l-o. 이드 099 2 2,-브 = 2-33 2 -33x o.8333 0.25 1-9441589 0-25 2 F.76356 60 31G 28S mext dgit